# BiasLab: Toward Explainable Political Bias Detection with Dual-Axis Annotations and Rationale Indicators

### v1

@@ Abstract
We present BiasLab, a dataset of 300 political news articles annotated for perceived ideological bias. These articles were selected from a curated 900-document pool covering diverse political events and source biases. Each article is labeled by crowdworkers along two independent scales, assessing sentiment toward the Democratic and Republican parties, and enriched with rationale indicators. The annotation pipeline incorporates targeted worker qualification and was refined through pilot-phase analysis. We quantify inter-annotator agreement, analyze misalignment with source-level outlet bias, and organize the resulting labels into interpretable subsets. Additionally, we simulate annotation using schema-constrained GPT-4o, enabling direct comparison to human labels and revealing mirrored asymmetries, especially in misclassifying subtly right-leaning content. We define two modeling tasks: perception drift prediction and rationale type classification, and report baseline performance to illustrate the challenge of explainable bias detection. BiasLab’s rich rationale annotations provide actionable interpretations that facilitate explainable modeling of political bias, supporting the development of transparent, socially aware NLP systems. We release the dataset, annotation schema, and modeling code to encourage research on human-in-the-loop interpretability and the evaluation of explanation effectiveness in real-world settings. Code & data: https://github.com/ksolaiman/PoliticalBiasCorpus
Index Terms:
political bias, media framing, perception annotation, LLM evaluation, crowdsourcing, dataset
I Introduction
@@
Political news reporting often reflects ideological leanings not just through overt opinion, but also through subtle cues such as selective emphasis, framing, or omission. While such framing is not inherently deceptive, it can reinforce partisan divides, obscure nuance, and influence public understanding of policy issues [1, 2]. Detecting such bias is not only a classification problem but also an explainability challenge: practitioners need models that can surface why a story reads as left- or right-leaning, not just that it does. While computational approaches to bias detection have made notable progress, most rely on coarse-grained supervision such as outlet-level bias labels [3, 4] or focus narrowly on framing in issue-specific corpora [5, 6]. Such proxies fail to capture how individual readers perceive bias within the same article.
@@
Subjective interpretation is central to this task. Prior work has shown that perceptions of media slant vary with reader background, political orientation, and article structure [11, 8]. Yet few datasets in NLP directly model this perception, especially at the document level, with full context and justification. Some efforts annotate sentence-level bias [6] or analyze framing in limited topics [5], while others simulate reader opinion using weak supervision [8]. However, there remains a lack of high-quality, article-level resources that capture perceived bias, account for inter-annotator disagreement, and support explainability through rationale [12, 13].
@@
We introduce BiasLab, a dataset of 300 U.S. political news articles annotated for perceived ideological slant. Articles are sampled from a curated pool of 900 partisan news stories covering diverse political events. Each article is labeled by crowdworkers along two independent Likert scales: one for sentiment toward the Democratic Party, and another toward the Republican Party. In addition to these dual-axis labels, annotators select rationale indicators adapted from media literacy frameworks [12, 13] and highlight relevant snippets. These rationales serve as built-in explanations, enabling downstream models not only to detect bias but also to justify their predictions. A qualification test and pilot-phase refinement ensure annotation quality [10].
@@
To benchmark the task beyond human annotation, we simulate labeling using GPT-4o under the same schema. This enables direct comparison of human and model judgments, revealing consistent asymmetries in perceived bias, most notably, a shared tendency to interpret subtly right-leaning content as neutral or even left-leaning. While recent work probes LLM–human alignment across moral reasoning [14], preference-guided evaluation [15], and grammaticality judgments [16], this work extends this line by examining alignment and misalignment of LLMs and humans in the domain of political news bias detection.
@@
BiasLab bridges an important gap between source-level ideology and document-level perception. It enables fine-grained study of ideological framing, disagreement modeling, and LLM alignment. We release the dataset, schema, baseline code, and resolved subsets to support future work on socially grounded NLP, bias detection and media understanding.
II Related Work
@@
Detecting political bias in news media has attracted considerable interest across computational linguistics, journalism studies, and political science. Existing research spans manual assessments at the source level to detailed analyses, including sentence-level annotations and event-specific framing studies. Prominent manual rating systems, such as the AllSides Media Bias Ratings [3] and the Ad Fontes Media Bias Chart [4], provide widely-used outlet-level ideological classifications but lack granularity at the article or sentence levels.
@@
Computational methods have expanded bias detection efforts to more detailed annotations within news texts. The Media Frames Corpus [5] annotates framing techniques related to specific political issues but does not directly address ideological bias labeling. Hamborg et al.’s WCL corpus [6] offers sentence-level bias annotations; however, its scale and granularity restrict broader applicability at the document level.
@@
Recently Machine learning approaches have significantly advanced political ideology detection. Iyyer et al. [7] utilized recursive neural networks to predict political ideology based on congressional speeches, a distinct context from general news media. Roy et al. [8] investigated framing identification in immigration discussions on social media, but their approach did not explicitly provide supervised labels at the article level.
@@
Crowdsourcing has emerged as an effective methodology for subjective annotation tasks, including bias detection. Budak et al. [11] demonstrated the effectiveness of crowd workers in quantifying media bias, emphasizing scalability and inter-annotator reliability. Nonetheless, challenges related to annotator agreement and managing subjective perceptions remain prevalent [10, 11]. Recent work has begun to scrutinize LLM–human alignment across several dimensions. Garcia et al. introduce a Moral Turing Test, revealing systematic gaps between human moral choices and LLM responses [14]. Liu et al. demonstrate that incorporating pairwise human-preference feedback markedly improves an evaluator model’s agreement with human bias judgments [15]. Complementing these findings, Hu et al. report strong model–human alignment on key grammatical constructions, suggesting that alignment quality can vary sharply by linguistic phenomenon [16]. Closer to our application domain, Baheti et al. analyze stance and safety interventions in offensive open-domain dialogue, highlighting practical challenges for bias-aware text generation [17]. While these studies chart important territory in alignment, document-level political perception remains under-explored; our work fills this gap by benchmarking schema-constrained GPT-4o labels against human dual-axis bias annotations in real-world news.
@@
III Dataset Construction
III-A Article Collection and Source Bias Assignment
@@
III-B Curated Subset and MTurk Task Preparation
From this large collection, we manually curated a dataset of 900 political news articles drawn from 414 distinct events. The selection criteria included balance of topic coverage, known standing of the outlets, and narrative clarity. Events included contentious legislative debates, political rallies, protests, and high-profile political developments between 2016–2018. Each article was linked to its source event and assigned an outlet-level bias using third-party assessments from AllSides [3] and Ad Fontes Media [4], identifying each as Left, Lean Left, Lean Right, or Right. While the full dataset includes articles from centrist outlets, the 900 curated subset selected for annotation contains only clearly partisan sources (i.e., Left or Right aligned). We then selected 300 articles from this curated pool for crowd-sourced annotation via Amazon Mechanical Turk (MTurk). This final set forms the core of the BiasLab dataset and is released alongside metadata for the broader 900-article pool to support future annotation efforts and replication studies. Among the 300 articles, 202 came from outlets considered Right-aligned or Leaning Right and 98 of them came from outlets considered Left-aligned or Leaning Left.
IV Annotation Methodology
IV-A Crowdsourced Annotation Task Design
@@
To support interpretability and explainability, a third question asked annotators to explain their ratings by selecting from a list of predefined bias indicators adapted from media literacy frameworks [12, 13]:
@@
Annotators could select multiple reasons and were instructed to highlight relevant text in the snippets to justify their selections. Annotators could optionally provide a brief rationale for their selections. To support detection of ‘omission bias’, we provided a short, neutral event description and summaries of both parties’ core stances.
Figure 2: Partial View of MTurk Bias Identification Task
IV-B Annotator Qualification and Quality Assurance
To ensure label quality, we restricted participation to U.S.-based workers with at least 1,000 approved HITs and an approval rating above 96%. Before accessing the main task, workers were required to pass a qualification test designed to assess their ability to recognize different forms of ideological bias. Use of qualifications tests to remove spammers and low quality workers is a common and effective practice on Mechanical Turk [10].
The test included seven multiple-choice questions that presented excerpts from news articles and asked workers to identify types of bias present. Each question included detailed feedback explaining the correct answer, serving both as a training step and a filter to ensure baseline annotation competency. As shown in figure 3, we added two questions for testing the capability of identifying ‘Bias by Word Choice and Tune’.
Figure 3: ‘Bias by word choice and tone’ in MTurk Qualification Test
@@ IV-B Annotator Qualification and Quality Assurance
IV-C Lessons from Pilot Phase
@@
V Annotation Evaluation and Agreement Analysis
@@
V-A Label Aggregation and Conflict Resolution Strategy
@@
[…]
V-B Annotator-Outlet Agreement and Conflict Analysis
To better understand how perceived bias aligns with the source outlet’s known political leaning, we conducted a post-hoc analysis of the annotations. We categorized the results into four groups, as summarized in Table I:
TABLE I: Summary of human annotation agreement versus outlet bias
@@ V-B Annotator-Outlet Agreement and Conflict Analysis
V-C Quantitative Alignment with Outlet Bias
@@
TABLE II: Per-Class Performance (Human Label vs. Outlet Bias)
@@
Table II summarizes per-class performance metrics for the same 270 articles. While source bias serves only as a weak supervisory signal, the class-level scores illustrate both alignment and divergence trends. Annotators performed most consistently on articles from right-leaning outlets (F1 = 0.58), likely due to more overt partisan cues in those articles. In contrast, left-leaning articles yielded more balanced precision and recall, while the Center class appears with zero scores.
@@
V-D Comparison with GPT-4o Annotations
@@
  • The model’s most common error was labeling Right articles as Left. This pattern is also observed among human annotators, but humans lean towards Center more.
@@
V-E Final Gold Set
We defined three subsets for supervised learning and evaluation:
@@ V-E Final Gold Set
VI Modeling Experiments
@@
VI-A Perception Drift Identification
@@
As shown in Table III, the model achieved an overall accuracy of 55.6%, with notably higher recall for match cases than drift cases. The class imbalance and linguistic subtlety of drift examples likely contribute to the lower performance on mismatches. These results validate the difficulty of detecting perceived bias drift from surface lexical features alone and establish a benchmark for future work.
TABLE III: Logistic Regression performance on perception drift identification (TF-IDF)
@@
VI-B Bias Rationale Type Classification
@@
We train one-vs-rest Logistic Regression models for each rationale type using TF-IDF features. Table IV summarizes macro-averaged results for the most common rationale types. The full per-class breakdown is shown in Appendix Table V.
TABLE IV: Macro-averaged Logistic Regression performance for rationale type classification.
@@ VII Discussion and Limitations
First, we observe that annotator agreement is strongest when bias cues are overt, and diverges when bias is conveyed through framing, omission, or emotional tone. This subjectivity reflects the nature of the task, and not a flaw. Capturing disagreement is essential for modeling real-world perception. BiasLab’s dual-axis design and rationale indicators help structure this ambiguity, enabling both explainability and fine-grained evaluation.
@@
We include initial modeling experiments to benchmark dataset utility and illustrate the difficulty of perception-driven classification. These pilots use explainable, replicable classifiers, and we leave deeper model tuning for future work.

### v2

# BiasLab: Toward Explainable Political Bias Detection with Dual-Axis Human Annotations and Rationale Indicators

@@ Abstract
We present BiasLab, a dataset of 300 political news articles annotated for perceived ideological bias. These articles were selected from a curated 900-document pool covering diverse political events and source biases. Each article is labeled by crowdworkers along two independent scales, assessing sentiment toward the Democratic and Republican parties, and enriched with rationale indicators adapted from media literacy guidelines. The annotation pipeline incorporates targeted worker qualification and was refined through pilot-phase analysis. We quantify inter-annotator agreement, analyze misalignment with source-level outlet bias, and organize the resulting labels into interpretable subsets. We complement human annotations with schema-constrained large language model (LLM) labeling to compare and align human and machine interpretability. These annotations captured mirrored asymmetries, especially in misclassifying subtly right-leaning content. We define two modeling tasks: perception drift prediction and rationale type classification, and report baseline performance to illustrate the challenge of explainable bias detection. Our analysis reveals notable disagreement patterns and perception drift, underscoring the subjectivity inherent in bias perception. We release the dataset, annotation schema, and baseline code, enabling feedback-driven alignment research using structured human labels and rationales for political bias perception.
Code & data: https://github.com/ksolaiman/PoliticalBiasCorpus
political bias detection, explainable AI, interpretable NLP, dataset annotation, dual-axis sentiment, rationale annotations, human annotation, large language models, perception, crowdsourcing
1 Introduction
@@
Political news reporting often reflects ideological leanings not just through overt opinion, but also through subtle cues such as selective emphasis, framing, or omission. While such framing is not inherently deceptive, it can reinforce partisan divides, obscure nuance, and influence public understanding of policy issues (1; 2). Detecting such bias is not only a classification problem but also an explainability challenge: practitioners need models that can surface why a story reads as left- or right-leaning, not just that it does. While computational approaches to bias detection have made notable progress, most rely on coarse-grained supervision such as outlet-level bias labels (3; 4) or focus narrowly on framing in issue-specific corpora (5; 6). Such proxies fail to capture how individual readers perceive bias within the same article.
@@
Subjective interpretation is central to this task. Prior work has shown that perceptions of media slant vary with reader background, political orientation, and article structure (11; 8). Yet few datasets in NLP directly model this perception, especially at the document level, with full context and justification. Some efforts annotate sentence-level bias (6) or analyze framing in limited topics (5), while others simulate reader opinion using weak supervision (8). However, there remains a lack of high-quality, article-level resources that capture perceived bias, account for inter-annotator disagreement, and support explainability through rationale (12; 13).
@@
We introduce BiasLab, a dataset of 300 U.S. political news articles annotated for perceived ideological slant. Articles are sampled from a curated pool of 900 partisan news stories covering diverse political events. Each article is labeled by crowdworkers along two independent Likert scales: one for sentiment toward the Democratic Party, and another toward the Republican Party. In addition to these dual-axis labels, annotators select rationale indicators adapted from media literacy frameworks (12; 13) and highlight relevant snippets. These rationales serve as built-in explanations, enabling downstream models not only to detect bias but also to justify their predictions. A qualification test and pilot-phase refinement ensure annotation quality (10). Table 1 shows a concrete example illustrating annotation schema and rationale tags.
Example Annotation Entry Title: Anti-Trump celebs plan ’People’s State of the Union’ event Event: President Trump will deliver his first State of the Union Article Snippet (excerpt): A group of Hollywood elites, progressive groups, and other Trump opponents are planning a “People’s State of the Union” to counter the president’s first official address. The event, coordinated by unions, organizers of the Women’s March and Planned Parenthood, is being marketed as a celebration of the “resistance,” closer to “the people’s point of view,” USA Today reported. Marked Bias Indicators: Marginalization of one side (Indicator 4): “A group of Hollywood elites … celebration of the resistance” Emotionally charged language (Indicator 0): “Hollywood elites,” “social activists,” “public alternative” Worker Labels: Right, Right Final Human Label: Right Outlet Bias: Right
Table 1: An example article from the dataset showing final labels, annotator agreement, and highlighted rationale types.
@@
To benchmark the task beyond human annotation, we simulate labeling using GPT-4o under the same schema. This enables direct comparison of human and model judgments, revealing consistent asymmetries in perceived bias, most notably, a shared tendency to interpret subtly right-leaning content as neutral or even left-leaning. While recent work probes LLM–human alignment across moral reasoning (14), preference-guided evaluation (15), and grammaticality judgments (16), BiasLab complements these by modeling alignment and misalignment of LLMs and human feedback in the domain of political news.
@@
BiasLab offers more than raw annotations: its structured schema transforms political bias perception into a testbed for human–model alignment. By capturing both directional sentiment and annotated rationale, it enables models to learn not just what bias humans perceive, but why. This positions BiasLab as a feedback-rich benchmark for training and evaluating alignment methods that go beyond label replication, grounded in public perception rather than rigid source-level ideology.
2 Related Work
@@
Detecting political bias in news media has attracted considerable interest across computational linguistics, journalism studies, and political science. Existing research spans manual assessments at the source level to detailed analyses, including sentence-level annotations and event-specific framing studies. Prominent manual rating systems, such as the AllSides Media Bias Ratings (3) and the Ad Fontes Media Bias Chart (4), provide widely-used outlet-level ideological classifications but lack granularity at the article or sentence levels.
@@
Computational methods have expanded bias detection efforts to more detailed annotations within news texts. The Media Frames Corpus (5) annotates framing techniques related to specific political issues but does not directly address ideological bias labeling. Hamborg et al.’s WCL corpus (6) offers sentence-level bias annotations; however, its scale and granularity restrict broader applicability at the document level.
@@
Recently Machine learning approaches have significantly advanced political ideology detection. Iyyer et al. (7) utilized recursive neural networks to predict political ideology based on congressional speeches, a distinct context from general news media. Roy et al. (8) investigated framing identification in immigration discussions on social media, but their approach did not explicitly provide supervised labels at the article level.
@@
Crowdsourcing has emerged as an effective methodology for subjective annotation tasks, including bias detection. Budak et al. (11) demonstrated the effectiveness of crowd workers in quantifying media bias, emphasizing scalability and inter-annotator reliability. Nonetheless, challenges related to annotator agreement and managing subjective perceptions remain prevalent (10; 11). Recent work has begun to scrutinize LLM–human alignment across several dimensions. Garcia et al. introduce a Moral Turing Test, revealing systematic gaps between human moral choices and LLM responses (14). Liu et al. demonstrate that incorporating pairwise human-preference feedback markedly improves an evaluator model’s agreement with human bias judgments (15). Complementing these findings, Hu et al. report strong model–human alignment on key grammatical constructions, suggesting that alignment quality can vary sharply by linguistic phenomenon (16). Closer to our application domain, Baheti et al. analyze stance and safety interventions in offensive open-domain dialogue, highlighting practical challenges for bias-aware text generation (17). While these studies chart important territory in alignment, document-level political perception remains under-explored; our work fills this gap by benchmarking schema-constrained GPT-4o labels against human dual-axis bias annotations in real-world news.
While scalar preference signals, such as pairwise comparisons (15) are widely used in alignment pipelines, recent work has explored more structured forms of human feedback, including multi‐choice critiques (19) and preference rankings (18). BiasLab complements these approaches by providing fine-grained rationales, label-reason mapping, flexible multi‐dimensional ratings, and explanatory feedback, and can serve as interpretable supervision for alignment tasks in politically sensitive domains.
@@
3 Dataset Construction
3.1 Article Collection and Source Bias Assignment
@@
3.2 Curated Subset and MTurk Task Preparation
From this large collection, we manually curated a dataset of 900 political news articles drawn from 414 distinct events. The selection criteria included balance of topic coverage, known standing of the outlets, and narrative clarity. Events included contentious legislative debates, political rallies, protests, and high-profile political developments between 2016–2018. Each article was linked to its source event and assigned an outlet-level bias using third-party assessments from AllSides (3) and Ad Fontes Media (4), identifying each as Left, Lean Left, Lean Right, or Right. While the full dataset includes articles from centrist outlets, the 900 curated subset selected for annotation contains only clearly partisan sources (i.e., Left or Right aligned). We then selected 300 articles from this curated pool for crowd-sourced annotation via Amazon Mechanical Turk (MTurk). This final set forms the core of the BiasLab dataset and is released alongside metadata for the broader 900-article pool to support future annotation efforts and replication studies. Among the 300 articles, 202 came from outlets considered Right-aligned or Leaning Right and 98 of them came from outlets considered Left-aligned or Leaning Left.
4 Annotation Protocol and Feedback Schema
4.1 Crowdsourced Annotation Task Design
@@
To support interpretability and explainability, a third question asked annotators to explain their ratings by selecting from a list of predefined bias indicators adapted from media literacy frameworks (12; 13):
@@
The seven rationale types align with three overarching dimensions of media bias: language and framing, balance and fairness, and factual integrity, as summarized in Appendix Figure 8. Annotators could select multiple reasons and were instructed to highlight relevant text in the snippets to justify their selections. Annotators could optionally provide a brief rationale for their selections. Table 1 presents a representative annotated article with marked bias indicators and final labels. To support detection of ‘omission bias’, we provided a short, neutral event description and summaries of both parties’ core stances.
Figure 2: Bias Identification Task
4.2 Annotator Qualification and Quality Assurance
To ensure label quality, we restricted participation to U.S.-based workers with at least 1,000 approved HITs and an approval rating above 96%. Before accessing the main task, workers were required to pass a qualification test designed to assess their ability to recognize different forms of ideological bias. Use of qualifications tests to remove spammers and low quality workers is a common and effective practice on Mechanical Turk (10).
The test included seven multiple-choice questions that presented excerpts from news articles and asked workers to identify types of bias present. Each question included detailed feedback explaining the correct answer, serving both as a training step and a filter to ensure baseline annotation competency. As shown in figure 3, we added two questions for testing the capability of identifying ‘Bias by Word Choice and Tone’.
Figure 3: Qualification Test Questions
@@ 4.2 Annotator Qualification and Quality Assurance
4.3 Lessons from Pilot Phase
@@
5 Feedback Quality and Human–Model Alignment
@@
5.1 Label Aggregation and Conflict Resolution Strategy
@@
This choice reflects two practical considerations: (1) center labels in such cases often reflect uncertainty rather than neutrality, and (2) at least one worker showed a pattern of overusing ‘Center’, likely as a safe default. To preserve specificity and reduce noise, we prioritized the more directional label in these cases. Our gold dataset also includes a final partisan score which is an average of the likert scale numbers from the two questions in the task. Negative values for the final partisan score indicated a left bias and positive values indicated a right bias for the article. Finally, a score with $0\pm 0.5$ was considered centrist.
5.2 Annotator-Outlet Agreement and Conflict Analysis
To better understand how perceived bias aligns with the source outlet’s known political leaning, we conducted a post-hoc analysis of the annotations. We categorized the results into four groups, as summarized in Table 2:
Table 2: Summary of human annotation agreement versus outlet bias
@@ 5.2 Annotator-Outlet Agreement and Conflict Analysis
5.3 Quantitative Alignment with Outlet Bias
@@
Table 3: Per-Class Performance (Human Label vs. Outlet Bias)
@@
Table 3 summarizes per-class performance metrics for the same 270 articles. While source bias serves only as a weak supervisory signal, the class-level scores illustrate both alignment and divergence trends. Annotators performed most consistently on articles from right-leaning outlets (F1 = 0.58), likely due to more overt partisan cues in those articles. In contrast, left-leaning articles yielded more balanced precision and recall, while the Center class appears with zero scores.
@@
5.4 Comparison with GPT-4o Annotations
@@
  • The model’s most common error was labeling Right articles as Left, which is also observed among human annotators, but humans lean towards Center more.
@@
5.5 Final Gold Set
We define a gold-standard subset of 270 articles derived from MTurk annotations, based on strict agreement criteria between independent annotators. This set is used for supervised learning, evaluation, and perception analysis:
@@ 5.5 Final Gold Set
Auxiliary Labels.
In addition to the human-annotated gold set, we release two supplementary label sources:
  • Outlet-derived bias labels for all 300 annotated articles and the broader 900-document corpus, based on the political leanings of their news sources. These serve as weak supervision signals and enable source-aware modeling.
  • Schema-constrained GPT-4o annotations for all 300 articles, designed to replicate the MTurk labeling scheme. While not part of the human gold set, these enable LLM–human agreement analysis and consistency benchmarking.
Together, these provide a flexible foundation for both perception-based learning and benchmark modeling across human, source, and machine perspectives.
6 Feedback-Based Benchmark Models
@@
6.1 Perception Drift Identification
@@
As shown in Table 4, the model achieved an overall accuracy of 55.6%, with notably higher recall for match cases than drift cases. The class imbalance and linguistic subtlety of drift examples likely contribute to the lower performance on mismatches. These results validate the difficulty of detecting perceived bias drift from surface lexical features alone and establish a benchmark for future work.
Table 4: Logistic Regression performance on perception drift identification (TF-IDF)
@@
6.2 Bias Rationale Type Classification
@@
We train one-vs-rest Logistic Regression models for each rationale type using TF-IDF features. Table 5 summarizes macro-averaged results for the most common rationale types. The full per-class breakdown is shown in Appendix Table 6.
Table 5: Macro-averaged Logistic Regression performance for rationale type classification.
@@ 7 Discussion and Future Work
First, we observe that annotator agreement is strongest when bias cues are overt, and diverges when bias is conveyed through framing, omission, or emotional tone. This subjectivity reflects the nature of the task, and not a flaw. Capturing disagreement is essential for modeling real-world perception. BiasLab’s dual-axis design and rationale indicators help structure this ambiguity, enabling both interpretability and fine-grained evaluation.
@@
We include initial modeling experiments to benchmark dataset utility and illustrate the difficulty of perception-driven classification. These pilots use interpretable, replicable classifiers, and we leave deeper model tuning for future work. Together, BiasLab offers a replicable, perception-grounded platform for evaluating media bias, interpretability, and alignment between human and machine.
8 Conclusion
BiasLab introduces a new resource for explainable political bias detection, pairing dual-axis sentiment labels with rationale indicators that surface why an article reads as left- or right-leaning. Our human-vs-LLM comparisons expose where both groups falter, especially on subtle framings, and our two baseline tasks (perception drift and rationale classification) underline the real-world difficulty of building transparent bias models. Moving forward, we plan to integrate richer contextual signals and interactive, human-in-the-loop workflows to strengthen explanation quality. We invite the community to explore new modeling approaches that treat structured human judgments as actionable alignment signals for perception-driven language model behavior.
9 Limitations